OpenStudy (anonymous):

A rain gutter is to be constructed of aluminum sheets 12 inches wide. After marking off a length of 4 inches from each edge, this length is bent up at an angle Θ. The area A of the opening may be expressed as the function: A(Θ) = 16 sin Θ • (cos Θ + 1). If Θ = 45°, what is the area of the opening?

OpenStudy (anonymous):

4.8 in2 9.0 in2 19.3 in2 20.8 in2

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@PeterPan

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@TheCatMan @phi @PeterPan @campbell_st

OpenStudy (phi):

They did not make you find the equation ... they told you it. Now use it. A(Θ) = 16 sin Θ • (cos Θ + 1) with Θ = 45° (which is one of the angles where people memorize the sin , cos and tan of) See http://www.allmathwords.org/en/e/exactvaltrig.html for a short table. Use those values in your equation. can you do that ?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i dont get it

OpenStudy (phi):

although, based on the decimal answers, you can just use a calculator

OpenStudy (anonymous):

what would i put in the calculator?

OpenStudy (phi):

are you saying you can't read A(Θ) = 16 sin Θ • (cos Θ + 1) ? 16 means 16 usually we do not bother to put in a multiplication sign, and 16 sin Θ means 16 times sin Θ in this case Θ is 45 degrees so you do 16 times sin(45) what do you get (from your calculator) ?

OpenStudy (anonymous):

i dont know how to input the sin

OpenStudy (anonymous):

or the cos

OpenStudy (anonymous):

i have a a scientific calculator

OpenStudy (phi):

why don't you take a few minutes and practice. type sin 30 enter what do you get for an answer?

OpenStudy (phi):

or sin 30 =

OpenStudy (anonymous):

.5

OpenStudy (phi):

ok, good. sin(30º)= ½ = 0.5 now do sin 45 =

OpenStudy (anonymous):

.707

OpenStudy (phi):

yes. so you can do sin now let's do the problem A(Θ) = 16 sin Θ • (cos Θ + 1) let's do the stuff inside the parens first do cos 45 =

OpenStudy (anonymous):

cos 45 is the same thing mate

OpenStudy (phi):

yes, sin45 = cos45 now do + 1 = it should add 1 to 0.707 to give 1.707

OpenStudy (anonymous):

k i did it

OpenStudy (phi):

if it says 1.707 you are good to continue type in x sin 45

OpenStudy (anonymous):

my calculator is in degrees mode

OpenStudy (anonymous):

thats how its supposed to be right?

OpenStudy (phi):

yes, I know it is in degree mode (they gave us theta= 45º the circle means degrees)

OpenStudy (anonymous):

i got 46

OpenStudy (anonymous):

is that right?

OpenStudy (phi):

we lost the bubble somewheres. start over and be careful. cos 45 + 1 = what do you get ?

OpenStudy (anonymous):

1.707

OpenStudy (phi):

now type x sin 45 x 16 =

OpenStudy (anonymous):

there is no x key do you want me to multiply?

OpenStudy (phi):

yes, whatever the multiply key is

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

sorry im used to seeing a dot for multipliation

OpenStudy (phi):

I was looking at my calculator and it has an "x"

OpenStudy (anonymous):

19.31

OpenStudy (anonymous):

im using the windows calculator app

OpenStudy (anonymous):

so some symbols are different

OpenStudy (phi):

yes. if your calculator has parens you can type in 16 * sin 45 * ( cos 45 + 1) = in one line.

OpenStudy (anonymous):

is 19.31 the right answer?

OpenStudy (phi):

yes, that looks good

OpenStudy (anonymous):

thanks could you help with another?

OpenStudy (phi):

ok

OpenStudy (anonymous):

What sine function represents an amplitude of 1, a period of 2π, a horizontal shift of π, and a vertical shift of −4? f(x) = sin (x − π) − 4 f(x) = sin (πx) + 4 f(x) = sin (x + pi over 2) + 4 f(x) = sin (pi over 2x) − 4

OpenStudy (anonymous):

how would i start this off?

OpenStudy (phi):

you use the "generic" equation \[ A \sin\left( \frac{2 \pi}{T} x - B\right)+C \] A is the amplitude T is the period B is the horizontal shift C is the vertical shift they told you A, T, B and C put them in, simplify where possible, and what do you get ?

OpenStudy (anonymous):

k let me solve

OpenStudy (anonymous):

wait what do i do for the x

OpenStudy (anonymous):

@phi

OpenStudy (phi):

the x stays as is

OpenStudy (anonymous):

k thanks man im sorry for bugging you

OpenStudy (phi):

any progress?

OpenStudy (anonymous):

f(x) = sin (x − π) − 4 is this the answer @phi

OpenStudy (phi):

yes

OpenStudy (anonymous):

k one more or no?

OpenStudy (anonymous):

because your a great help man